Timberly Construction negotiates a lump-sum purchase of several assets from a company that is going out of business. The purchase is completed on January 1, 2017, at a total cash price of $810,000 for a building, land, land improvements, and four vehicles. The estimated market values of the assets are building, $446,200; land, $320,100; land improvements, $48,500; and four vehicles, $155,200. The company’s fiscal year ends on December 31.


Required:

1-a. Prepare a table to allocate the lump-sum purchase price to the separate assets purchased.
1-b. Prepare the journal entry to record the purchase.
2. Compute the depreciation expense for year 2017 on the building using the straight-line method, assuming a 15-year life and a $32,000 salvage value.
3. Compute the depreciation expense for year 2017 on the land improvements assuming a five-year life and double-declining-balance depreciation.
Answer #1
1a
Allocation of total cost Appraised Value Percent of Total
Appraised Value
x Total cost of
Acquisition
Apportioned Cost
Building 446,200 46% x 810,000 372,600
Land 320,100 33% x 810,000 267,300
Land improvements 48,500 5% x 810,000 40,500
Vehicles 155,200 16% x 810,000 129,600
Total 970,000 100% 810,000
b
Date General Journal Debit Credit
1-Jan Building 372,600
Land 267,300
Land improvements 40,500
Vehicles 129,600
Cash 810,000
2
Depreciation expense on building 22707 =(372600-32000)/15
3
Depreciation rate 40% =1/5*2
Depreciation expense on
land improvements
16200 =40500*40%